myininaya (myininaya):

\[\text{ Show } x^2y^2(4-x^2)(4-y^2) \le 16 \] \[\text{ for all real numbers } x \text{ and } y \text{ such that } \\ |x| \le 2 \text{ and } |y| \le 2\]

OpenStudy (fibonaccichick666):

I don't buy that for all x,y , is there a restriction?

myininaya (myininaya):

changed the question to reflect the changes

OpenStudy (fibonaccichick666):

kk lol

OpenStudy (fibonaccichick666):

so if \[|x|\le2\] then \[|x|^2=|x^2|=x^2\le2^2\]

OpenStudy (fibonaccichick666):

similar for y, yes?

myininaya (myininaya):

ok

OpenStudy (fibonaccichick666):

eh wait, maybe just use the difference of squares, what topic are you on in class atm?

myininaya (myininaya):

Well can talk about any approach we want.

myininaya (myininaya):

But I can tell you I got it from my special book of magical calculus spells. (It is just james stewart calculus 6th edition) I have one approach in mind but like to see what you guys can come up with as well. Like any way you guys can think of.

myininaya (myininaya):

Even ways that would hurt my brain.

OpenStudy (fibonaccichick666):

For the above, I was going to use a limit as it approaches x,y |4| So second idea \[x^2y^2(4−x^2)(4−y^2)≤16\] could become: \[x^2y^2(2−x)(2+x)(2−y)(2+y)≤16\] so then we can say the maximum of each is obtained at 0 or 2, so we end up with a zero in the multiplication string when each individual attains it's max. which discredits those

OpenStudy (fibonaccichick666):

but the smart thing would probably be to apply a derivative and determine that way

OpenStudy (fibonaccichick666):

we would have to show that the derivative was never positive

OpenStudy (fibonaccichick666):

we would have to use partial derivatives in a R3 environment, I think

myininaya (myininaya):

\[ \text{ Let } f(z)=z^2(4-z^2) \text{ where we have } -2 \le z \le 2 \] \[f(\pm 2)=0\] \[f'(z)=4z(2-z^2)\] \[f'=0 \text{ gives } z=0 \text{ or } z=\pm \sqrt{2}\] \[\text{ but } f(0) \text{ doesn't yield our max value }\] \[f(\pm \sqrt{2})=4 \text{ which is the max value }\] \[ \text{ So } f(z) \le 4 \] \[f(x) \cdot f(y) \le 4 \cdot 4 =16\]

myininaya (myininaya):

Is the way I was thinking

myininaya (myininaya):

but i want to see the partials too

OpenStudy (fibonaccichick666):

I don't recommend using z, make it something else there. but is that the derivative? let me see...

OpenStudy (fibonaccichick666):

also, based on your given, you would have that expression squared

OpenStudy (fibonaccichick666):

but we don't know that x=y

OpenStudy (fibonaccichick666):

i feel like we would need to set it up like this in 3 space(maybe, \(\color\orange{\text{IE if you know how to do this just substitute yourself into the game }}\))

OpenStudy (fibonaccichick666):

no... i can't think of how to parameterize it... hmmm

myininaya (myininaya):

"I don't recommend using z, make it something else there. but is that the derivative? let me see... " why is this? "also, based on your given, you would have that expression squared " what expression? - also x and y do not have to be equal but f(x) and f(y) will still both have max value 4

myininaya (myininaya):

on the intervals from x=-2 to x=2 and from y=-2 to y=2 that is

OpenStudy (fibonaccichick666):

so partials, would yield: \[x^2y^2(4−x^2)(4−y^2)≤16\\ f_x=y^2(4-y^2)(2x(4-x^2)+x^2*-2x)\\ f_y=x^2(4-x^2)(2y(4-y^2)+y^2*-2y)\\f_z=0\]

OpenStudy (fibonaccichick666):

and z implies complex plane alot(a heck of a lot for me lately) so I try not to use it for real numbers anymore so I don't comfuse myself on roots

OpenStudy (fibonaccichick666):

the expression was your change of variable to Z, I assumed you let x and y be the same in which case your z expression would have been squared

OpenStudy (fibonaccichick666):

what about switching to polar?

myininaya (myininaya):

I didn't really need to introduce another variable I could have found the max values for f(x)=x^2(4-x^2) on x<= 2 and f(y)=y^2(4-y^2) on y<= 2 but I thought doing the same thing twice was a waste of time

myininaya (myininaya):

|x|<=2 * |y|<=2*

OpenStudy (fibonaccichick666):

yea... hmmm

OpenStudy (anonymous):

hmm x^2y^2<=16 (4-y^2)(4-x^2)<=16 but the relation btw those two terms is inverse xD like maximize x^2y^2 ,means minimize (4-y^2)(4-x^2) ...ect we need some cool theorem hmm

OpenStudy (fibonaccichick666):

yea, I mean they cancel each other out essentially

OpenStudy (anonymous):

they dont cancel ...

OpenStudy (anonymous):

somehow , we need to consider cases less than 16 :)

ganeshie8 (ganeshie8):

*

myininaya (myininaya):

you know you guys consider the max values on the intervals from -2 to 2 for x^2 and (4-x^2) separately but why can't we consider the max value of the product of x^2 and (4-x^2)

OpenStudy (fibonaccichick666):

how do we find that max value, optimization

myininaya (myininaya):

well that is what I did in my post above

OpenStudy (fibonaccichick666):

i feel like the max should be 3

OpenStudy (anonymous):

ok here is something since bounds are equivalent i'll consider |x|=|y| x^4(4-x^2)^2 and for less values assume x^4=s x^4(16-8x^2+x^4) 16x^4 -8x^6+x^8 s^2-8s^(3/2)+16s

OpenStudy (fibonaccichick666):

well, 9

OpenStudy (anonymous):

eh i'll see now if that works

myininaya (myininaya):

\[f(x)=x^2(4-x^2) \text{ on } x \in [-2,2] \\ f'(x)=2x(4-x^2)+x^2(-2x)=-4x^3+8x=-4x(x^2-2) \\ \\ f'=0 \text{ gives } x=0 \text{ or } x=\pm \sqrt{2} \\ \text{ so plug in our criticqal numbers along with our endpoints into } f\]

myininaya (myininaya):

to see what would be our max value

myininaya (myininaya):

f(-2)=0 f(2)=0 f(0)=0 f(-sqrt(2))=4 f(sqrt(2)=4

myininaya (myininaya):

our max value is 4

myininaya (myininaya):

for f(x)=x^2(4-x^2)

myininaya (myininaya):

so our max value for f(y)=y^2(4-y^2) is 4

OpenStudy (fibonaccichick666):

ok so at sqrt of 2 yea

OpenStudy (fibonaccichick666):

I'll be back later to check, I have an assignment due at midnight

OpenStudy (anonymous):

thats it ? xD

myininaya (myininaya):

well i mean i'm interested in other ways if there are other ways

OpenStudy (anonymous):

there is L M theorem but im too tired to remember it xD

myininaya (myininaya):

L M?

OpenStudy (anonymous):

but then i thought its only polynomial

myininaya (myininaya):

lagrange multiplier ?

OpenStudy (anonymous):

yes let me see ifi found my books

OpenStudy (anonymous):

wait not yes lagrange sorry

OpenStudy (anonymous):

yes when u said L M , so i replied yes

ganeshie8 (ganeshie8):

that product of functions thingy is a neat trick @myininaya :) finding critical points isn't that complicated here as everything is factored already

OpenStudy (anonymous):

ok cant find it anywhere XD

ganeshie8 (ganeshie8):

\[f(x,y)=x^2y^2(4-x^2)(4-y^2)\] \(f_x = y^2(4-y^2)(8x-4x^3)\) we dont need to work the other partial as we can use symmetry \(f_x = 0 \implies y = 0 , \pm 2\), \(x = 0, ~x = \pm \sqrt{2}\) the function feels continuous every where in the given domain so it has to achieve maximum value at critical points or the boundary points

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

why do you want to use lagrange multipliers here ?

ganeshie8 (ganeshie8):

oh we can interpret the given bounds as constraints is it ?

OpenStudy (anonymous):

no one wanna use it

myininaya (myininaya):

do you guys just want to do another problem? or do you guys have fun problems you can post?

OpenStudy (anonymous):

if u have post

myininaya (myininaya):

ok I will just choose a random one

ganeshie8 (ganeshie8):

Alright, lagrange multipiliers makes no sense to me here.. .i admit im confused a bit about seeing the difference between "bounds" and "constraint function"

ganeshie8 (ganeshie8):

il stick to the modest way of finding critical points or the other two methods that were discussed in this thred :)

OpenStudy (anonymous):

wat modest ?

OpenStudy (anonymous):

what other two way

ganeshie8 (ganeshie8):

1) myininaya's method of taking product of max values of two functions of different variables : http://gyazo.com/eca9604182d11a282ad7f7e235a4f898 2) your method which you know :) 3) finding critical values

ganeshie8 (ganeshie8):

i consider last method is modest because thats the method textbooks teach and also thats the method you turn to when your brain stops working :P